1.25x – 2y = 1.3125
x + 2y = –0.75

Answer by Alan3354(69443)

1.25x – 2y = 1.3125
x + 2y = –0.75
-------------
Multiply by 16 to get rid of fractions. Not required, but it makes it easier.
20x - 32y = 21
16x + 32y = -12
Add 2 eqns
36x = 9
x = 0.25
-------
Sub in 2nd eqn
.25 + 2y = -.75
2y = -1
y = -0.5